# Towards Foundation Models for Experimental Readout Systems Combining Discrete and Continuous Data

### v1

@@ 5 Summary and Conclusions
With this work, we have introduced a (proto) Foundation Model working on low-level input features (i.e., pixelated readout systems) for Nuclear Physics applications. We have shown through a series of closure tests that our model is capable of performing both generative, and reconstruction tasks in which we are able to leverage fine-tuning to increase computational efficiency.
@@
In the case of generative modeling, our discretized approach in space allows potential for capturing low-level geometric effects seen in DIRC detectors, an advancement over prior methods [15]. Moreover, we have introduced a method in which we alleviate the need for excess quantization over continuous domains, allowing the physical sensor resolutions of detectors to be respected through adjacent vocabularies. Our method fuses spatial and temporal information through Causal Multi-Head Cross Attention (CMHCA), yielding temporally aware spatial embeddings that are subsequently processed using standard transformer blocks. We believe this architecture is both flexible and broadly applicable to a variety of detector systems, including calorimeters. Given the inherent stochasticity in physical readout systems, the concept of a many-to-many mapping approximated through split vocabularies may offer promising results.
We have also demonstrated that our model achieves the desired separation power ( at 6 GeV/c) for charged pion and kaon identification, with performance comparable to or exceeding that of other Deep Learning methods. Interestingly, our Foundation Model exhibits greater sensitivity to variations in polar angle, which may reflect its internal representation of detector geometry or the nature of its kinematic conditioning. Additionally, we find that fine-tuning not only accelerates convergence but also yields slight gains in classification accuracy. Notably, the model demonstrates excessively strong discriminative power at lower momenta—though the practical importance of such performance may be limited. To better align model focus with experimental priorities, future studies will explore momentum weighted, or limited phase-space training strategies that emphasize high-impact kinematic regions.
@@
Despite the encouraging performance observed, further refinement of our method is necessary to enable higher-fidelity simulation and improved classification performance, particularly at higher momentum. As seen in Large Language Models, emergent capabilities often arise as a function of scale. Accordingly, we intend to investigate the scalability of our approach through detailed ablation studies focused on both model capacity and the volume of training data.
@@
Lastly, we aim to assess the model’s ability to fine-tune across experiments, with particular interest in the GlueX DIRC and the hpDIRC at the future EIC. While differences in production mechanisms and detector geometries exist between GlueX and the EIC, we hypothesize that the underlying physics governing DIRC detectors remains sufficiently invariant to enable effective cross-experiment fine-tuning.
@@ Acknowledgments
We thank William & Mary for supporting the work of JG and CF through CF’s start-up funding. The authors acknowledge William & Mary Research Computing for providing computational resources and technical support that have contributed to the results reported within this article.

### v2

@@ 5 Summary and Conclusions
We present a (proto) foundation model designed to process low-level detector inputs (e.g., pixelated readout systems), with demonstrated use in nuclear physics contexts and potential relevance to other particle detection environments. Through a series of closure tests, we have demonstrated that our model can perform both generative and reconstruction objectives, leveraging fine-tuning to improve computational efficiency for targeted applications.
@@
In the case of generative modeling, our discretized approach in space allows potential for capturing low-level geometric effects seen in DIRC detectors, an advancement over prior methods [18]. Moreover, we have introduced a method in which we potentially alleviate the need for excess quantization over continuous domains, allowing the physical sensor resolutions of detectors to be respected through adjacent vocabularies. Our method fuses spatial and temporal information through Causal Multi-Head Cross Attention, yielding temporally aware spatial embeddings that are subsequently processed using standard transformer blocks. We further show that multiple classes (e.g., pions and kaons) can be contained within a singular generative model using a Mixture of Experts without loss of granularity in comparison to multiple independent models. This is a key advancement over previous methods in which independent models were used to prevent over mixing of modes as momentum increases—crucial for particle identification applications where the probability density function differs marginally at the hit level. We believe this architecture is both flexible and broadly applicable to a variety of detector systems, including calorimeters. Given the inherent stochasticity in physical readout systems, the concept of a many-to-many mapping approximated through split vocabularies may offer promising results. We have also demonstrated that our model achieves the desired separation power ( at 6 GeV/c) for charged pion and kaon identification, with performance comparable to or exceeding that of other Deep Learning methods [17]. Interestingly, our foundation model exhibits enhanced sensitivity to variations in polar angle, potentially reflecting its internal representation of detector geometry or the influence of kinematic conditioning—an effect that will be investigated in future work. Additionally, we find that fine-tuning not only accelerates convergence but also yields slight gains in classification accuracy. Notably, the model demonstrates exceptionally high distinguishing power at lower momenta—though the practical importance of such performance may be limited. To better align model focus with experimental priorities, future studies will explore momentum weighted, or limited phase-space training strategies that emphasize high-impact kinematic regions.
Finally, we have demonstrated that our architecture can also be applied to noise filtering—a critical task for high-noise sub-detector systems at the future EIC-via token-level classification. In this context, fine-tuning offers limited benefit, as the model must effectively “unlearn” global context in favor of localized information. While the difference in performance between a model trained from scratch and one that is fine-tuned is marginal, we have identified potential limitations of fine-tunability within reconstruction tasks.
@@
Despite the encouraging performance observed, further refinement of our method is necessary to enable higher-fidelity simulation and improved classification performance, particularly at higher momentum. As seen in Large Language Models, emergent capabilities often arise as a function of scale [34]. Accordingly, we intend to investigate the scalability of our approach through detailed ablation studies focused on both model capacity and the volume of training data.
@@
Lastly, we aim to assess the model’s ability to fine-tune across experiments, with particular interest in the GlueX DIRC [35] and the hpDIRC at the future EIC [36]. While differences in production mechanisms and detector geometries exist between GlueX and the EIC, we hypothesize that the underlying physics governing DIRC detectors remains sufficiently invariant to enable effective cross-experiment fine-tuning.
@@ Acknowledgments
The authors acknowledge William & Mary Research Computing for providing computational resources and technical support that have contributed to the results reported within this article.